# Unveiling the nature of SN 2022jli: the first double-peaked stripped-envelope supernova showing periodic undulations and dust emission at late times

### v1

@@
We present optical and infrared observations from maximum light until around days of supernova (SN) 2022jli, a peculiar stripped envelope (SE) SN showing two maxima, each one with a peak luminosity of about erg s-1 and separated by fifty days. The second maximum is followed by unprecedented periodic undulations with a period of days. The spectra and the photometric evolution of the first maximum are consistent with the behaviour of a standard SE SN with an ejecta mass of , and a radioactive 56Ni mass of . The optical spectra after days relative to the first maximum correspond to a standard SN Ic event, and at late times SN 2022jli exhibits a significant drop in the optical luminosity implying that the physical phenomena that produced the secondary maximum has ceased to power the SN light curve. One possibility is that the secondary maximum is powered by a magnetar with an initial spin period of ms and a high magnetic field of G, while the light curve periodic undulations could be produced by accretion of material from a companion star onto the neutron star in a binary system. The near-infrared spectra shows clear 1st CO overtone emission from about days after the first maximum, and it becomes undetected at days. A significant near-infrared excess from hot dust emission is detected at days produced by either newly formed dust in the SN ejecta or due to a strong near-infrared dust echo. Depending on the assumptions of the dust composition, the estimated dust mass is . The magnetar power of the second maximum can fit in a more general picture where magnetars are the power source of super-luminous SNe, could produce their frequent bumps and undulations, and where pulsars could produce the late time excess observed in some SE SNe. The detection of CO and the potential detection of dust formed in the ejecta of SN 2022jli are important to understand the formation molecules and dust in the ejecta of SE SNe.
@@
Massive stars with zero-age main-sequence (ZAMS) masses larger than end their lives as core-collapse supernovae (SNe), giving birth to neutron stars and black holes. These stellar explosions can be broadly divided in hydrogen-rich Type II SNe exhibiting strong and long-lasting hydrogen features in their spectra, and hydrogen-deficient stripped-envelope (SE) SNe which have lost the majority of their hydrogen and helium envelopes over their evolutionary lifetimes (see Filippenko, 1997; Gal-Yam, 2017).
@@
A scenario initially proposed for SE SNe was the explosion of single massive Wolf-Rayet (WR) stars that have lost their envelopes through strong winds (Conti, 1975). In this scenario, SE SNe (IIb, Ib and Ic) follow a sequence in mass, where the most massive stars suffer from stronger winds loosing a larger fraction of their envelopes before demise as a SN event. This predicts that the most massive stars demise as SNe Ic loosing all of their hydrogen and most, if not all, of their helium envelopes.
@@
An alternative progenitor scenario for SE SNe are massive stars in binary systems that shed much of their envelopes through binary interaction (Podsiadlowski et al., 1992), while in the late-stage of their evolution, stellar winds can play a key role in peeling off the remaining hydrogen (Ib’s) and helium (Ic’s) in their outermost layers (see e.g., Yoon, 2017; Dessart et al., 2020). This scenario allows a larger range of initial stellar masses to become SE SNe, but still the most massive stars lose a larger fraction of their envelopes and explode as SNe Ic (Yoon, 2017; Dessart et al., 2020). This scenario is supported by the large fraction of young massive stars in binary systems (Sana et al., 2012), and by their small ejecta masses () ranging between about (Drout et al., 2011; Cano, 2013; Taddia et al., 2015; Lyman et al., 2016; Prentice et al., 2016; Taddia et al., 2018b), which are too small to be consistent with single massive stars. Their radioactive 56Ni mass () is in the range of from sample studies of SE SNe (Drout et al., 2011; Cano, 2013; Taddia et al., 2015; Lyman et al., 2016; Prentice et al., 2016; Taddia et al., 2018b; Anderson, 2019).
@@
Standard SE SNe display bell-shaped light curves having a single maximum (see e.g., Taddia et al., 2015, 2018a),111Some SE SNe exhibit an early bump a few days after the explosion and several days before the 56Ni powered maximum (eg., SN 1993J; see Richmond et al., 1994). This early excess emission is explained by shock cooling of extended material. powered by the decay of radioactive 56Ni. SN 2005bf (Monard et al., 2005) was the first SE SN reported to show a peculiar light curve morphology composed of two maxima separated by about 25 days (Anupama et al., 2005; Tominaga et al., 2005; Folatelli et al., 2006; Maeda et al., 2007; Stritzinger et al., 2018b). The canonical radioactive decay of 56Ni distributed close to the SN center is not able to produce bumpy or double-peaked light curves. However, a double-peaked distribution of radioactive 56Ni, with an enhanced 56Ni abundance near the SN surface, provides a reasonable alternative to produce double-peaked light curves (see Orellana & Bersten, 2022, for a discussion). This was the first scenario proposed to explain the light curves of SN 2005bf (Tominaga et al., 2005; Folatelli et al., 2006), however it faces the difficulty to reconcile the relatively high 56Ni mass predicted with the fast decline and the faint SN luminosity at about a year after the explosion (see Maeda et al., 2007). Motivated by this discrepancy, Maeda et al. (2007) proposed that the energy injection from a newly born magnetar could explain the maximum light morphology, the rapid decline and the faint luminosity of SN 2005bf at late times. Since then, PTF11mnb (Taddia et al., 2018a) and SN 2019cad (Gutiérrez et al., 2021) are two SE SNe reported to display similar light curves morphology to SN 2005bf. In both cases their light curves can be modelled with a double-peaked distribution of 56Ni. Gutiérrez et al. (2021) show that the light curves of SN 2019cad can be equally modelled with a central distribution of 56Ni and power injection from a magnetar, which powers the second maximum. Another noteworthy event is SN 2022xxf (Itagaki, 2022). This SN was discovered before maximum and promptly classified as a SN Ic broad line (Ic-BL; Balcon, 2022). It shows two prominent maxima separated by days, the second maximum reached a slightly brighter luminosity and bluer colours, compared with the first maximum (Kuncarayakti et al., 2023). The first maximum seems to correspond to a standard SN Ic-BL event, while the second maximum has been argued to be powered by the interaction between the SN ejecta with a massive (a few ) hydrogen and helium poor circum-stellar medium (CSM; see Kuncarayakti et al., 2023). This hypothesis is based on spectral features suggesting ejecta-CSM interaction and on the appearance of a conspicuous intermediate width line component ( km s-1) on top of the broad emission ( km s-1). The spectral evolution is slow and after the second maximum the light curve declines rapidly.
@@
Hydrogen-deficient super-luminous SE SNe (SLSNe-Ic) correspond to a luminous class of objects encompassing peak luminosities from erg s-1 or mag (see e.g., De Cia et al., 2018; Lunnan et al., 2018; Angus et al., 2019) to erg s-1 or mag (see e.g., De Cia et al., 2018; Lunnan et al., 2018; Angus et al., 2019; Cartier et al., 2022). Their large peak luminosities are difficult to produce through standard SN powering mechanisms such as the canonical decay of 56Ni synthesized in the SN explosion. The most popular scenarios to explain the light curves of SLSNe-Ic are the power injection from a magnetar (Kasen & Bildsten, 2010; Woosley, 2010; Inserra et al., 2013), or strong interaction with hydrogen and helium poor CSM (see e.g., Chatzopoulos et al., 2013; Sorokina et al., 2016; Moriya et al., 2018, for a discussion). Pre-maximum bumps are often found in the light curves of the SLSNe-Ic population (Leloudas et al., 2012; Nicholl et al., 2015; Smith et al., 2016; Nicholl & Smartt, 2016; Angus et al., 2019). These early bumps are usually explained by the shock cooling emission from extended material surrounding the progenitor star (see Nicholl & Smartt, 2016; Piro, 2015; Piro et al., 2021). Subsequent bumps, humps and light curve undulations are frequently found throughout SLSNe-Ic light curve evolution (Yan et al., 2015; Nicholl et al., 2016b; Yan et al., 2017; Hosseinzadeh et al., 2022; Cartier et al., 2022; Gutiérrez et al., 2022). Some bumps are likely explained by the interaction between the SN ejecta with a dense CSM (Yan et al., 2015, 2017), while in other cases they can be explained by the variable power injection from a magnetar (see Moriya et al., 2022).
@@
A combination of mechanisms such as the decay of radioactive nickel, power injection from a magnetar and interaction between the SN ejecta with a dense CSM are invoked to explain the double-peaked light curve of the luminous SN 2019stc (Gomez et al., 2021). Early bumps can also be the result of the collision between the SN ejecta and a companion star, yielding double-peaked light curves (Kasen, 2010), the result of magnetar driven shock-break out (Kasen et al., 2016), or the collision of the newborn neutron star with the companion star in a binary system (see Hirai & Podsiadlowski, 2022).
@@
SN 2022jli is a SE SN discovered near maximum light in the nearby galaxy NGC 157. SN 2022jli shows a strong secondary maximum followed by periodic or quasi-periodic undulations (Moore et al., 2023; Chen et al., 2024). Chen et al. (2024) proposed that the first maximum corresponds to a SE SN, and the second maximum and the subsequent light curve undulations are produced by the accretion from a companion star onto a compact object at periastron encounters in a binary system (see also King & Lasota, 2024). Alternatively, Moore et al. (2023) proposed that the second maximum corresponds to a massive SN ejecta powered by the radioactive decay of 56Ni, and the periodic variability could a consequence of ejecta-CSM interaction with the CSM configured in nested shells or due to the collision of a neutron star with the companion star as in the model of Hirai & Podsiadlowski (2022).
This work is structured as follows. In Section 2 we present optical and near-infrared (NIR) observations of SN 2022jli from maximum light to days. In Section 3 we present the analysis of the observations of SN 2022jli, we estimate the host galaxy reddening in the line-of-sight, we present and analyse the bolometric light curve of SN 2022jli. We show that the SN is characterized by two maxima of similar luminosity, where the first maximum corresponds to a standard SN Ic, we estimate the SN ejecta mass and the 56Ni mass under the assumption that the first maximum is powered by the radioactive decay of 56Ni. We confirm the undulations in the optical light curves that appear after the second maximum, finding a period of about 12.5 days, consistent with Moore et al. (2023) and Chen et al. (2024). We study the spectral sequence starting from days to days, showing that during the secondary maximum the optical and NIR spectral features evolve slowly, implying that the SN ejecta remains heated by more than days. At late times ( days) the SN suffers a drop in luminosity reaching a luminosity comparable to a standard SE SN at about days, and the optical spectra after days are typical of a SN Ic at nebular phase. We report the formation of carbon monoxide (CO) from the detection of the 1st CO overtone in the NIR spectra at days and the emission of hot dust at late times ( days) detected by means of a clear infrared excess. In Section 4, we discuss the super-Eddington accretion scenario (Chen et al., 2024; King & Lasota, 2024) and the SN ejecta-CSM interaction scenario (Moore et al., 2023), and propose an alternative scenario in which in addition to the accretion of material from a companion star onto the neutron star, the power from the spin-down rotational energy from a magnetar contribute to power the second maximum. This alternative scenario alleviates the need of an extreme beaming to explain the high luminosity observed during the second maximum, and can be placed within the broad picture of bumps and undulations observed in magnetar powered SLSNe and late time excess potentially produced y pulsars in some SE SNe (see e.g., Ravi et al., 2023). In Section 5 we summarise our conclusions.
@@ 2 Observations and Data Reduction
SN 2022jli was discovered on 2022 May 05.17 UT by Libert Monard (Monard, 2022) at an unfiltered (clear filter) apparent magnitude of 14.4 mag, which was later corrected to mag and reported to the Transient Name Server (TNS). The SN is located in the galaxy NGC 157 at equatorial coordinates of and (J2000), in the northern spiral arm of its host galaxy, at west and north of the galaxy centre (see Fig. 1). SN 2022jli was classified as a SN Ic at a phase of one to three days after maximum based on a spectrum obtained on 2022 May 11.14 UT by Grzegorzek (2022), and later confirmed as a SN Ic by ePESSTO+ (Cosentino et al., 2022) based on a spectrum obtained on 2022 May 24.42 UT. Due to some confusion on the SN coordinates reported to TNS, the ePESSTO+ classification was erroneously reported as a separate object named SN 2022jzy.
@@ 2.1 Photometry
We analyse public optical photometry from the Kleinkaroo Observatory reported by Monard to the TNS, from the Gaia mission222http://gsaweb.ast.cam.ac.uk/alerts/alert/Gaia22cbu/ (Gaia Collaboration et al., 2016), from the ATLAS survey (Tonry et al., 2018; Smith et al., 2020), from the Zwicky Transient Facility (ZTF; Bellm et al., 2019), and from All-Sky Automated Survey for Supernovae (ASAS-SN; Shappee et al., 2014). We also present optical photometry obtained with the LDSS-3 instrument mounted to the 6.5-m Magellan Clay telescope, the EFOSC2 instrument with the 3.58-m New Technology Telescope (NTT), the Goodman instrument mounted on the 4.1 m SOAR telescope, and NIR photometry obtained with the Flamingos-2 instrument on the 8.4-m Gemini-South observatory. The optical photometry extends from a few days before maximum light to days after maximum, and the NIR photometry extends from days to days.
@@ 2.1.1 Optical
To discount any potential issue in the public ZTF photometry of SN 2022jli we downloaded the original ZTF images from the NASA/IPAC Infrared Science Archive,333https://irsa.ipac.caltech.edu/Missions/ztf.html and performed host-galaxy template image subtraction using HOTPANTS (Becker, 2015). Aperture and PSF photometry was then computed from these template subtracted science images using Python scripts. ATLAS forced photometry from the ATLAS forced photometry server was also downloaded and included in our analysis.444https://fallingstar-data.com/forcedphot/ We include -band ASAN-SN photometry to complement the photometry during the decline from the first maximum and on the rise to the second maximum. The ASAN-SN photometry correspond to photometry published by Moore et al. (2023) and Chen et al. (2024).
The reduction of the LDSS-3 and Goodman images was performed in IRAF following standard procedures that include bias subtraction and the normalization of the science images using a normalized flat image, the astrometric registration of the images was done using astrometry.net (Lang et al., 2010). The reduction of the EFOSC2 -band acquisition image was done with the PESSTO pipeline (Smartt et al., 2015). Aperture and PSF photometry was performed with Python scripts (see e.g., Hueichapan et al., 2022). The optical photometry of SN 2022jli is presented in Fig. 2.

### v2

@@
We present optical and infrared observations from maximum light until around days of supernova (SN) 2022jli, a peculiar stripped-envelope (SE) SN showing two maxima, each one with a peak luminosity of about erg s-1, separated by 50 days. The second maximum is followed by unprecedented periodic undulations with a period of days. The spectra and the photometric evolution of the first maximum are consistent with the behaviour of a standard SE SN with an ejecta mass of and a radioactive 56Ni mass of . The optical spectra after days relative to the first maximum correspond to a standard SN Ic event, and at late times SN 2022jli exhibits a significant drop in the optical luminosity, implying that the physical phenomena that produced the secondary maximum have ceased to power the SN light curve. Among other potential scenarios, we discuss how the second maximum could be powered by a magnetar, while the light curve periodic undulations could be produced by accretion of material from a companion star onto the neutron star in a binary system. The near-infrared spectra shows clear first CO overtone emission from about days after the first maximum, and it becomes undetected at days. A significant near-infrared excess from hot dust emission is detected at days, having been produced by either newly formed dust in the SN ejecta or a strong near-infrared dust echo. Depending on the assumptions of the dust composition, the estimated dust mass is . The potential magnetar power of the second maximum can fit into a more general picture in which magnetars are the power source of SE super-luminous SNe, and could explain bumps, undulations, and late-time excess emission in SE SNe. The CO detection and the dust emission of SN 2022jli are key to understanding the molecule and dust formation in the ejecta of SE SNe and in their environment.
@@
Massive stars with a zero-age main-sequence (ZAMS) mass higher than end their lives as core-collapse supernovae (SNe), giving birth to neutron stars and black holes. These stellar explosions can be broadly divided into hydrogen-rich Type II SNe exhibiting strong and long-lasting hydrogen features in their spectra, and hydrogen-deficient stripped-envelope (SE) SNe that have lost the majority of their hydrogen and helium envelopes over their evolutionary lifetimes (see Filippenko, 1997; Gal-Yam, 2017).
@@
An initial scenario proposed for SE SNe was the explosion of single massive Wolf-Rayet (WR) stars that have lost their envelopes through strong winds (Conti, 1975). In this scenario, SE SNe (IIb, Ib, and Ic) follow a sequence in mass, whereby the most massive stars suffer from stronger winds, losing a larger fraction of their envelopes before their demise as a SN event. This predicts that the most massive stars demise as SNe Ic, losing all of their hydrogen and most, if not all, of their helium envelopes.
@@
An alternative progenitor scenario for SE SNe is massive stars in binary systems that shed much of their envelopes through binary interaction (Podsiadlowski et al., 1992). At the same time, in the late stages of their evolution, stellar winds can play a key role in peeling off the remaining hydrogen (Ib’s) and helium (Ic’s) in their outermost layers (see e.g., Yoon, 2017; Dessart et al., 2020). This scenario allows a larger range of initial stellar masses to become SE SNe, but still, the most massive stars lose a larger fraction of their envelopes and explode as SNe Ic (Yoon, 2017; Dessart et al., 2020). This scenario is supported by the large fraction of young massive stars in binary systems (Sana et al., 2012), and by their small ejecta masses () ranging between about (Drout et al., 2011; Cano, 2013; Taddia et al., 2015; Lyman et al., 2016; Prentice et al., 2016; Taddia et al., 2018b), which are too small to be consistent with single massive stars. Their radioactive 56Ni mass () is in the range of from sample studies of SE SNe (Drout et al., 2011; Cano, 2013; Taddia et al., 2015; Lyman et al., 2016; Prentice et al., 2016; Taddia et al., 2018b; Anderson, 2019; Sharon & Kushnir, 2020; Afsariardchi et al., 2021; Rodríguez et al., 2023).
@@
Standard SE SNe display bell-shaped light curves that have a single maximum (see e.g., Taddia et al., 2015, 2018a),111Some SE SNe exhibit an early bump a few days after the explosion and several days before the 56Ni powered maximum (eg., SN 1993J; see Richmond et al., 1994). This early excess emission is explained by shock cooling of extended material. powered by the decay of radioactive 56Ni. SN 2005bf (Monard et al., 2005) was the first SE SN reported to show a peculiar light curve morphology composed of two maxima separated by about 25 days (Anupama et al., 2005; Tominaga et al., 2005; Folatelli et al., 2006; Maeda et al., 2007; Stritzinger et al., 2018b). The canonical radioactive decay of 56Ni distributed close to the SN centre is not able to produce bumpy or double-peaked light curves. However, a double-peaked distribution of radioactive 56Ni, with an enhanced 56Ni abundance near the SN surface, provides a reasonable alternative to produce double-peaked light curves (see Orellana & Bersten, 2022, for a discussion). This was the first scenario proposed to explain the light curves of SN 2005bf (Tominaga et al., 2005; Folatelli et al., 2006); however, it faces the difficulty of reconciling the relatively high 56Ni mass predicted with the fast decline and the faint SN luminosity at about a year after the explosion (see Maeda et al., 2007). Motivated by this discrepancy, Maeda et al. (2007) proposed that the energy injection from a newly born magnetar could explain the maximum light morphology, the rapid decline, and the faint luminosity of SN 2005bf at late times. PTF11mnb (Taddia et al., 2018a) and SN 2019cad (Gutiérrez et al., 2021) are two SE SNe that display a similar light curve morphology to SN 2005bf. In both cases their light curves can be modelled with a double-peaked distribution of 56Ni; however, Gutiérrez et al. (2021) show that the light curves of SN 2019cad can equally be modelled with a central distribution of 56Ni and power injection from a magnetar, which produces the second maximum. Another noteworthy event is SN 2022xxf (Itagaki, 2022). This SN was discovered before maximum and promptly classified as a SN Ic broad line (Ic-BL; Balcon, 2022). It shows two prominent maxima separated by days. The second maximum reached a slightly brighter luminosity and bluer colours, compared with the first maximum (Kuncarayakti et al., 2023). The first maximum appears to correspond to a standard SN Ic-BL event. In turn, it has been argued that the second maximum was powered by the interaction between the SN ejecta and a massive (a few ) hydrogen- and helium-poor circumstellar medium (CSM; see Kuncarayakti et al., 2023). This hypothesis is based on spectral features suggesting ejecta-CSM interaction and on the appearance of a conspicuous intermediate-width line component ( km s-1) on top of the broad emission ( km s-1). The spectral evolution is slow, and after the second maximum the light curve declines rapidly. The hypothesis proposed by Kuncarayakti et al. (2023) requires further investigation to confirm such a massive hydrogen- and helium-poor CSM.
@@
Hydrogen-deficient super-luminous SE SNe (SLSNe-Ic) correspond to a luminous class of objects encompassing peak luminosities from erg s-1 or mag (see e.g., De Cia et al., 2018; Lunnan et al., 2018; Angus et al., 2019; Cartier et al., 2022) to erg s-1 or mag (see e.g., De Cia et al., 2018; Lunnan et al., 2018; Angus et al., 2019; Cartier et al., 2022). Their high peak luminosities are difficult to produce through standard SN powering mechanisms such as the canonical decay of 56Ni synthesised in the SN explosion. The most popular scenarios to explain the light curves of SLSNe-Ic are power injection from a magnetar (Kasen & Bildsten, 2010; Woosley, 2010; Inserra et al., 2013), or strong interaction with a hydrogen- and helium-poor CSM (see e.g., Chatzopoulos et al., 2013; Sorokina et al., 2016; Moriya et al., 2018, for a discussion). Pre-maximum bumps are often found in the light curves of the SLSNe-Ic population (Leloudas et al., 2012; Nicholl et al., 2015; Smith et al., 2016; Nicholl & Smartt, 2016; Angus et al., 2019). These early bumps are usually explained by the shock cooling emission from extended material surrounding the progenitor star (see Nicholl & Smartt, 2016; Piro, 2015; Piro et al., 2021). Subsequent bumps, humps, and light curve undulations are frequently found throughout SLSNe-Ic light curve evolution (Yan et al., 2015; Nicholl et al., 2016b; Yan et al., 2017; Hosseinzadeh et al., 2022; Cartier et al., 2022; Gutiérrez et al., 2022). Some bumps are likely explained by the interaction between the SN ejecta with a dense CSM (Yan et al., 2015, 2017), while in other cases they can be explained by the variable power injection from a magnetar (see Moriya et al., 2022).
@@
A combination of mechanisms such as decay of radioactive nickel, power injection from a magnetar, and interaction between the SN ejecta with a dense CSM are invoked to explain the double-peaked light curve of the luminous SN 2019stc (Gomez et al., 2021). Early bumps can also be the result of the collision between the SN ejecta and a companion star, yielding double-peaked light curves (Kasen, 2010), the result of magnetar-driven shock-break out (Kasen et al., 2016), or the collision of the newborn neutron star with the companion star in a binary system (see Hirai & Podsiadlowski, 2022).
@@
SN 2022jli is a SE SN discovered near maximum light in the nearby galaxy NGC 157. SN 2022jli shows a strong secondary maximum followed by periodic or quasi-periodic undulations (Moore et al., 2023; Chen et al., 2024). Chen et al. (2024) proposed that the first maximum corresponds to a SE SN, and the second maximum and the subsequent light curve undulations are produced by the accretion from a companion star onto a compact object at periastron encounters in a binary system (see also King & Lasota, 2024). Alternatively, Moore et al. (2023) proposed that the second maximum corresponds to a massive SN ejecta powered by the radioactive decay of 56Ni, and the periodic variability could a consequence of ejecta-CSM interaction with the CSM configured in nested shells or due to the collision of a neutron star with the companion star, as in the model of Hirai & Podsiadlowski (2022). During the revision of this article, accretion models from a shock-inflated companion star into a neutron star in a binary system appeared. These models seem to explain the second maximum and the periodic light curve undulations of SN 2022jli (Hirai et al., 2025; Lu et al., 2025). A recent hydrodynamical model that combines the radioactive decay of of 56Ni and a magnetar successfully reproduces the first and second maxima of SN 2022jli, but does not reproduce the periodic undulations of SN 2022jli (Orellana et al., 2025).
This work is structured as follows. In Section 2 we present optical and infrared (IR) observations of SN 2022jli from maximum light to days. In Section 3 we present the analysis of the observations of SN 2022jli. In Section 4, we discuss the super-Eddington accretion scenario (Chen et al., 2024; King & Lasota, 2024) and the SN ejecta-CSM interaction scenario (Moore et al., 2023), and propose an alternative scenario in which in addition to the accretion of material from a companion star onto the neutron star, the power from the spin-down rotational energy from a magnetar contributes to powering the second maximum. In Section 5 we summarise our conclusions.
@@ 2 Observations and data reduction
SN 2022jli was discovered on 2022 May 05.17 UTC by Libert Monard (Monard, 2022) at an unfiltered (clear filter) apparent magnitude of 14.4 mag, which was later corrected to mag and reported to the Transient Name Server (TNS). The SN is located in the galaxy NGC 157 at equatorial co-ordinates of :: and :: (J2000), in the northern spiral arm of its host galaxy, west and north of the galaxy centre (see Fig. 1). SN 2022jli was classified as a SN Ic at a phase of one to three days after maximum based on a spectrum obtained on 2022 May 11.14 UTC by Grzegorzek (2022), and later confirmed as a SN Ic by ePESSTO+ (Cosentino et al., 2022) based on a spectrum obtained on 2022 May 24.42 UTC. Due to some confusion about the SN co-ordinates reported to TNS, the ePESSTO+ classification was erroneously reported as a separate object named SN 2022jzy.
@@ 2.1 Photometry
We analysed public optical photometry from the Kleinkaroo Observatory reported by Monard to the TNS, from the Gaia mission222http://gsaweb.ast.cam.ac.uk/alerts/alert/Gaia22cbu/ (Gaia Collaboration et al., 2016), from the ATLAS survey (Tonry et al., 2018; Smith et al., 2020), from the Zwicky Transient Facility (ZTF; Bellm et al., 2019), and from All-Sky Automated Survey for Supernovae (ASAS-SN; Shappee et al., 2014). We also present optical photometry obtained with the LDSS-3 instrument mounted to the 6.5-m Magellan Clay telescope, the EFOSC2 instrument with the 3.58-m New Technology Telescope (NTT), the Goodman instrument mounted on the 4.1 m SOAR telescope, and NIR photometry obtained with the Flamingos-2 instrument on the 8.1-m Gemini-South observatory. The optical photometry extends from a few days before maximum light to days after maximum, and the NIR photometry extends from days to days.
@@ 2.1.1 Optical
To discount any potential issue in the public ZTF photometry of SN 2022jli, we downloaded the original ZTF images from the NASA/IPAC Infrared Science Archive,333https://irsa.ipac.caltech.edu/Missions/ztf.html and performed host-galaxy template image subtraction using HOTPANTS (Becker, 2015). Aperture and PSF photometry was then computed from these template subtracted science images using Python scripts. ATLAS forced photometry from the ATLAS forced photometry server was also downloaded and included in our analysis.444https://fallingstar-data.com/forcedphot/ We include -band ASAN-SN photometry to complement the photometry during the decline from the first maximum and on the rise to the second maximum. The ASAN-SN photometry correspond to photometry published by Moore et al. (2023) and Chen et al. (2024).
The reduction of the LDSS-3 and Goodman images was performed in IRAF following standard procedures that include bias subtraction and the normalisation of the science images using a normalised flat image, the astrometric registration of the images was done using astrometry.net (Lang et al., 2010). The reduction of the EFOSC2 -band acquisition image was done with the PESSTO pipeline (Smartt et al., 2015). Aperture and PSF photometry was performed with Python scripts (see e.g., Hueichapan et al., 2022). The optical photometry of SN 2022jli is presented in Fig. 2. LDSS-3, EFOSC2 and Goodman photometry is reported in Table 8 in the Appendix.